OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = -1/5 and point (5, -2). A. y + 2 = -(1/5)(x - 5) B. y + 5 = -(1/5)(x - 2) C. y – 5 = (1/5)(x + 2) D. y – 5 = (1/5)(x - 2)

OpenStudy (zehanz):

General formula for a line with slope m, through point (a,b) is y - b = m(x - a). So put in m=-1/5, a=5 and b = -2! Even before that, you can dismiss answers C and D...